A quiz has 8 questions with answer options A and B for each. What is the probability that by pure guess work a student answer at least 6 correctly?
Answer by stanbon(75887)

A quiz has 8 questions with answer options A and B for each. What is the probability that by pure guess work a student answer at least 6 correctly?
---------
It is a binomial problem with n=8, p=1/2, 6<=x<=8
---------
If you have a Ti calculator use 1- binomcdf(8,1/2,5)= 0.145
--------
If you do not have a TI you need to figure 8C6(1/2)^6(1/2)2+8C7(1/2)^7(1/2)
+8C8(1/2)8 = 37*(1/2)^8=0.145
==========
Cheers,
Stan H.